John Deere has an 7% coupon bond outstanding. It has a $1,000 face value, matures in exactly 14 years (28 semi-annual periods) and pays interest semi-annually. If current market rates for bonds of this quality are 5.4%, what is the bond's price?
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 14 * 2 = 28 semi-annual periods, i = 5.4%/2 = 2.7% (semi-annual), C = 7% * $1000/2 = $35 (semi-annually)

1 P = 35* (1+0.027)25 0.027 - + 1000 (1 + 0.027 28

P = $681.50 + $474.27

P = $1,155.77